In 1985 Maria was three times as old as Sue. Five years earlier, the sum of their ages was eighteen. In what year was Maria born?
Answer by Maths68(1474)


In 1985
========
Maria was = m years old
Sue was = s years old
Maria was three times as old as Sue
m=3s........(1)
1980
====
Five years earlier
(s-5)+(m-5)=18
s-5+m-5=18
s+m-10=18
s+m=18+10
s+m=28.......(2)
=============================================================================
Put the value of m from (1) to (2)
s+(3s)=28
s+3s=28
4s=28
4s/4=28/4
s=7
Put the value of s in (1)
m=3s
m=3(7)
m=21
In 1985
Maria was = m = 21 years old
Sue was = s = 7 = years old
Maira was born in = 1985 - 21 = 1964